8a.A house is advertised as having 1900 square feet under roof. What is the area of this house in square meters?


1
Expert's answer
2022-06-06T08:54:27-0400
"1\rm \: ft=0.3048\:\rm m"

"1\rm \: ft^2=(0.3048\:\rm m)^2=0.0929\:\rm m^2"

"1900\rm \: ft^2=1900*0.0929\:\rm m^2=176.5\: m^2"